Grade 10Trigonometry

If tanA= 1 and tan B=√ 3. Find cos A.cos B- sin A.sin B

Bhagyashree
6 Years agoGrade 10

1 Answer

Arun
6 Years ago
Dear student
 
tan (A + B) = 1 + sqrt3 / 1 – sqrt 3 
 
hence  cos (A + B) = (1 – sqrt 3 )/ 2sqrt2
 
also cosA cosB – sin A sinB = cos (A + B)